OpenStudy (anonymous):

Each dimension of a rectangular prism is increased by 10%. What percent of the original volume is the volume of the new prism? 00.1 percent 133.1 percent 121 percent 1 percent

OpenStudy (anonymous):

PLEAASE HELP ! thanks <3

OpenStudy (anonymous):

Let a,b,c be the lengths of the sides of the rectangular prism. Then we have: \[V_0=abc\] Increasing each by 10% we have: \[V_1=(1.1a)(1.1b)(1.1c)=1.1^3(abc)\] Using the ratio: \[\frac{V_1}{V_0}=\frac{1.1^3(abc)}{abc}=1.1^3=1.331\] Multiplying the ratio by 100 to get the percentage, we have \[\% = 1.331 * 100\% = 133.1\%\]

OpenStudy (anonymous):

so it is b ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thanks so muchh

OpenStudy (anonymous):

No problem!